# An efficient Monte Carlo method for valid prior-free possibilistic statistical inference

### v2

@@ 3.5 Back to a possibilistic IM
The starting point was a possibilistic IM with contour $\pi_{z}$ that generally can only be evaluated using Monte Carlo by sampling artificial data sets from the underlying statistical model. This is cheap and easy to do at a few parameter values, but becomes overwhelmingly expensive when the contour evaluations are needed at all parameter points on a sufficiently fine grid. Unfortunately, virtually every relevant summary of the IM output requires this evaluation on a fine grid. What was proposed above was a probabilistic approximation $\mathsf{Q}_{z}^{\star}$ of the possibilistic IM $\overline{\Pi}_{z}$, which could be sampled with relative ease. Then those relevant summaries, e.g., expectations, could be readily approximated using the samples from $\mathsf{Q}_{z}^{\star}$, and subsequently used to approximate the IM’s summaries. The challenge is that, for each summary, there are mathematical limits to how well the probabilistic approximation can match the target possibilistic one; for some summaries, the approximation is accurate while for others it’s not. The question here is if there’s a way to convert the probabilistic approximation $\mathsf{Q}_{z}^{\star}$ to a possibilistic approximation that better represents the target IM while retaining the computational efficiency.
The answer to this question is Yes. The “inner probabilistic approximation” just described can be interpreted as a possibility-to-probability transform, but a probability-to-possibility transform goes the other direction. Let $r_{z}:\mathbb{T}\to\mathbb{R}$ be a (possibly data-dependent) ranking function on the parameter space, where larger values of $r_{z}(\theta)$ indicate that $\theta$ is “higher ranked” in some meaningful sense. Good examples of ranking functions are, first, the density function $q_{z}^{\star}$ corresponding to the distribution $\mathsf{Q}_{z}^{\star}$ and, second, the likelihood function $L_{z}$ of the underlying model. With both of these choices, the meaningfulness of the ranking is clear. Taking the ranking function to be the density $q_{z}^{\star}$ gives the tightest contours, similar to how highest density sets have smallest volume among those with a fixed probability content, but “tightest contours” isn’t the objective. Moreover, there may be computational challenges associated with use of the density-based ranking. After all, $q_{z}^{\star}$ would need to be estimated using, say, kernel methods applied to the samples from $\mathsf{Q}_{z}^{\star}$ as described above, which is non-trivial in the multiparameter case. See below for more on how the ranking function affects the approximation accuracy.
@@
$\omega_{z}(\theta)=\mathsf{Q}_{z}^{\star}\bigl{\{}r_{z}(\Theta)\leq r_{z}(\theta)\bigr{\}},\quad\theta\in\mathbb{T}.$ (16)
@@
To see that $\sup_{\theta}\omega_{z}(\theta)=1$ and, hence, that this transformation defines a genuine possibility contour, just take $\theta$ equal (or converging) to a “highest ranked” value according to $r_{z}$. Given that $\omega_{z}$ is a genuine possibility contour, it makes sense to define the corresponding possibility measure via optimization as before, i.e., $\overline{\Omega}_{z}(H)=\sup_{\theta\in H}\omega_{z}(\theta)$. In the examples that follow, I’ll refer to the approximation $\omega_{z}$ of $\pi$ as the stitched IM contour, since it’s obtained by stitching together the $\alpha$-cut approximations.
@@
Computationally, the contour $\omega_{z}$ in (16) is far more efficient than that in (7). The reason being that the Monte Carlo samples in the former are fixed whereas, in the latter, the samples generally depend on the $\theta$ at which the contour is being evaluated. It’s also worth asking if $\omega_{z}$ is a good approximation of $\pi_{z}$. Towards this, if $\mathsf{Q}_{z}^{\star}$ is a genuine inner probabilistic approximation of $\overline{\Pi}_{z}$, i.e., if (10) holds exactly, then it’s easy to see that
$\mathsf{Q}_{z}^{\star}\{\pi_{z}(\Theta)\leq\pi_{z}(\theta)\}=\pi_{z}(\theta),\quad\theta\in\mathbb{T},$
where the left-hand side above is $\omega_{z}(\theta)$ corresponding to the ranking function $r_{z}=\pi_{z}$. So, with a proper choice of ranking function, it’s theoretically possible to recover the original IM from this. The downside is that, of course, if $\pi_{z}$ was available to serve as a ranking function, then there’d be no need for any of these approximations. So, it’s unrealistic to expect that the stitched IM with contour $\omega_{z}$ will exactly recover the original IM. That said, however, since the original IM is driven by the relative likelihood function, it follows that the $\pi_{z}$-ranking and the $L_{z}$-ranking are very similar. In fact, they are exactly the same if the relative likelihood $R(Z,\theta)$ is a pivot under $\mathsf{P}_{\theta}$, which is the case asymptotically for all regular models. To summarize, there is a sense in which the likelihood-based ranking is preferred, but this doesn’t mean the likelihood-based version of the stitched IM contour $\omega_{z}$ in (16) exactly matches that of the original IM. In the large-sample context discussed in Section 3.4 above, however, one can expect $\omega_{z}$ and $\pi_{z}$ to merge asymptotically, for both the likelihood- and density-based rankings.
4 Examples
The illustrative examples below are low-dimensional and, therefore, the naive approximation (7) of the IM contour is computationally feasible. So, my goal here is to offer some further computational details concerning the new proposal in Section 3 and to highlight its approximation quality, efficiency, etc. R code for the gamma illustration in Example 2 below is available at https://www4.stat.ncsu.edu/~rgmarti3/software; some more versatile and user-friendly software is currently in the works.
@@ 4 Examples
One of the more challenging inference problems involving a one-parameter model is the bivariate normal with known means and variances but unknown correlation. Suppose that $Z=(Z_{1},\ldots,Z_{n})$ are iid, with $Z_{i}=(X_{i},Y_{i})$ a bivariate normal random vector with zero means, unit standard deviations, and correlation $\Theta\in\mathbb{T}=(-1,1)$ to be inferred. Interestingly, that the means and variances are known makes the problem more difficult—it’s a curved exponential family so the minimal sufficient statistic is not complete and there are various ancillary statistics available to condition on (e.g., Basu, 1964). How this affects asymptotic inference is detailed in Reid, (2003). It is straightforward to construct a possibilistic IM (Martin, 2024a , Example 2), which is exactly valid for all sample sizes and asymptotically efficient, but computation of the naive approximation (7) is relatively expensive because there’s no pivotal structure and no closed-form expression for the maximum likelihood estimator. The variational approximation proposed in Cella and Martin, (2024) is fast and easy, but it loses the exact validity of the original IM. Here I apply the proposed Monte Carlo sampling strategy on Fisher’s transformation scale, i.e., on $\Psi=\text{arctanh}(\Theta)$, the inverse hyperbolic tangent. Figure 1(a) shows a histogram of 5000 samples from the distribution $\mathsf{Q}_{z}^{\star}$ of $\Psi$ as described above, based on (a centered and scaled version of) the law school admissions data analyzed in Efron, (1982), where $n=15$ and the maximum likelihood estimator of $\Theta$ is $\hat{\theta}_{z}=0.789$. Overlaid on this plot is, first, a normal density with mean $\text{arctanh}(0.789)=1.07$ and estimated standard deviation and, second, the kernel density estimate; the difference between these two estimates is negligible. Panel (b) shows the exact IM contour and three stitched IM approximations as in (16) corresponding to three different ranking functions $r_{z}$. All three approximations have roughly the same shape as the true contour, as expected. The two based on density estimates—Gaussian and kernel—are quite similar and closely agree with the exact contour. The likelihood-based approximation is very accurate on the right-hand side but a bit conservative on the left-hand side; this is because, as in Figure 1(a), the distribution $\mathsf{Q}_{z}^{\star}$ of the Fisher-transformed $\Theta$ is symmetric but the likelihood-based ranking is not.
(a) Samples from $\mathsf{Q}_{z}^{\star}$
(b) Possibility contours
Figure 1: Summary of the results in Example 1 for the bivariate normal correlation. In Panel (a), the solid line is the Gaussian density and dashed line is the kernel density estimate from R’s density. In Panel (b), the gray line is the exact IM contour, the dashed and dotted lines are approximations using the Gaussian and kernel density rankings, respectively, and the solid line is the approximation based on the likelihood ranking.
Example 2.
@@
Let $Z=(Z_{1},\ldots,Z_{n})$ denote an iid sample from a gamma distribution with unknown parameter $\Theta=(\Theta_{1},\Theta_{2})$, where $\Theta_{1}>0$ and $\Theta_{2}>0$ are the unknown shape and scale parameters, respectively. It is straightforward to get the naive approximation (7) at any particular value of the parameter, but prohibitively expensive to carry this out over a sufficiently fine grid that spans the plausible pairs $(\theta_{1},\theta_{2})$. So, there’s a need for more computationally efficient approximation methods, and the Monte Carlo strategy of Section 3 above fits the bill. For this illustration, I’ll work with the data presented in Example 3 of Fraser et al., (1997), which consists of the survival time (in weeks) for $n=20$ rats exposed to a certain amount of radiation, modeled by a gamma distribution with unknown shape and scale parameters. Figure 2 summarizes the 5000 samples of $(\log\Theta_{1},\log\Theta_{2})$ from the proposed probabilistic approximation $\mathsf{Q}_{z}^{\star}$ developed in Section 3. This reveals that the distribution is at least approximately Gaussian. Figure 3(a) shows the exact joint contour function for $(\Theta_{1},\Theta_{2})$ along with two stitched IM approximations: one takes the ranking function to be a bivariate Gaussian density function and the other takes it to be the gamma likelihood function. Notice that the likelihood-ranking-based approximation almost perfectly matches the exact contour. And for comparison, the exact contour requires more than 10 minutes of computation time, while the new Monte Carlo approximation is completed in roughly 10 seconds.
A practically relevant and surprisingly challenging follow-up question concerns inference on the mean $\Phi=\Theta_{1}\Theta_{2}$ of the gamma distribution. IM solutions that offer exactly valid inference on $\Phi$ are presented in Martin and Liu, 2015c and Martin, 2023b , but the relevant computations can be a burden. An obvious idea is that, given a sample of $(\log\Theta_{1},\log\Theta_{2})$ values from the probabilistic approximation, it’s straightforward to get a corresponding sample of $\Phi=\exp(\log\Theta_{1}+\log\Theta_{2})$ values, but would this provide a good approximation of the marginal IM for $\Phi$ presented in Martin, 2023b ? To answer this question, Figure 3(b) shows the latter exact, profile likelihood-based, marginal IM possibility contour for the mean $\Phi$ and two approximations thereof:
  • an indirect approximation based on the samples of $(\log\Theta_{1},\log\Theta_{2})$ from $\mathsf{Q}_{z}^{\star}$ as described above, using the profile likelihood as the ranking function, and
  • a direct approximation wherein the mean $\Phi$ is sampled from the inner probabilistic approximation of the target marginal IM.
From the context alone, the reader should expect that the direct approximation would be superior to the indirect approximation, but this superiority doesn’t come for free: the former is tailored to the particular choice of feature $\Phi$ and, therefore, is not as simple to obtain as the latter. From the plot in Figure 3(b), it’s clear that both approximations are quite accurate and, indeed, the direct approximation is more accurate overall, with the indirect approximation being a bit too narrow compared to the target. This narrowness is inevitable because the marginal distribution for $\Phi$ derived from $\mathsf{Q}_{z}^{\star}$ generally would not be the inner probabilistic approximation of the marginal IM. Whatever this approximation might lack in terms of accuracy, it makes up for in simplicity.
@@
Figure 2: Summary of the samples $(\log\Theta_{1},\log\Theta_{2})$ from the probabilistic approximation $\mathsf{Q}_{z}^{\star}$ of the IM in Example 2.
@@
where $F(u)=(1+e^{-u})^{-1}$ is the logistic distribution function. The corresponding likelihood cannot be maximized in closed-form, but this is easy to do numerically, and the maximum likelihood estimator and the corresponding observed information matrix lead to the asymptotically valid inference reported by standard statistical software. Figure 4 shows the data and the fitted “death probability” curve. For exact inference, however, the computational burden is heavier: evaluating the exact IM contour for $\Theta$ over a sufficiently fine grid of $\theta$ values again is prohibitively expensive. As an alternative, the Monte Carlo sampling method presented in Section 3 is easy to implement and runs in a matter of seconds. Figure 5(a) shows the 5000 Monte Carlo samples of $(\Theta_{1},\Theta_{2})$ from $\mathsf{Q}_{z}^{\star}$ along with the stitched IM contour based on a Gaussian density ranking, closely agreeing with the approximations in Cella and Martin, (2024) and Martin and Williams, (2024).
@@
Figure 4: Plot of the data described in Example 3 with the fitted “death probability” curve. The vertical line at $\hat{\lambda}_{z}=0.244$ marks the estimated median lethal dose (LD50), which is the exposure level at which the death probability is 0.5.
@@
Figure 5: Summary of the logistic regression results results in Example 3. Panel (a) shows the samples of $(\Theta_{1},\Theta_{2})$ from $\mathsf{Q}_{z}^{\star}$ and the corresponding approximate possibility contour. Panel (b) shows the approximate possibility contour for the median lethal dose $\Lambda$ based on the samples in Panel (a).
5 Application
@@
Following the review in Section 2, given that $\mathsf{G}$ is a nuisance parameter, the natural strategy is to work with a relative profile likelihood for $\Theta$ which, in this case, is given by
@@
$\hat{\pi}_{z}(\theta)=\mathsf{P}_{\theta,\widehat{G}}\bigl{\{}R^{\text{\sc pr}}(Z,\theta)\leq R^{\text{\sc pr}}(z,\theta)\bigr{\}},\quad\theta\in\mathbb{T}.$ (18)
Evaluation of the right-hand side via Monte Carlo boils down to sampling censoring levels from $\widehat{G}$, sampling concentration levels from $\mathsf{P}_{\theta}$, and then constructing new data sets according to (17). While this procedure is conceptually relatively simple, naive implementation over a sufficiently fine grid of $\theta$ values is very expensive. Fortunately, the proposed Monte Carlo strategy can be readily applied to sample from an inner probabilistic approximation of $\hat{\pi}_{z}$, from which approximately valid inference on $\Theta$ can be obtained. While only first-order asymptotic validity of this IM could be established in Cahoon and Martin, (2021), the empirical results presented there are quite striking in that they suggest the possibility of a higher-order accuracy similar to that associated with bootstrap.
@@
Figure 6: Plot of samples $(\Theta_{1},\Theta_{2})$ of the Weibull shape and scale parameters from the inner probabilistic approximation $\mathsf{Q}_{z}^{\star}$ along with two corresponding joint contours: Gaussian density ranking (solid) and the asymptotic Gaussian possibility contour (dashed) as developed in Martin and Williams, (2024).
While the Weibull model is common in applications, the Weibull model parameters themselves are difficult to interpret. A relevant and interpretable feature is the mean of the Weibull distribution, which is given by $\Phi=\Theta_{2}\,\Gamma(1+\Theta_{1}^{-1})$, where $\Gamma$ is the usual gamma function. Like in the gamma and logistic regression examples above, it is straightforward to obtain an approximate marginal IM for the relevant feature $\Phi$ or, equivalently $\log\Phi$, based on the samples of $(\Theta_{1},\Theta_{2})$ from $\mathsf{Q}_{z}^{\star}$. Figure 7(a) shows a histogram of the samples of $\log\Phi$ and, since there’s a slight sign of asymmetry, I use the kernel density estimate-based ranking function to construct the marginal contour for $\log\Phi$ in Figure 7(b). From here, a nominal 90% confidence interval for $\log\Phi$ can be immediately read off, i.e., $(6.41,7.74)$. Exponentiating the endpoints gives a corresponding nominal 90% confidence interval for the mean survival time $\Phi$ as $(610.7,2309.0)$.
@@
Figure 7: Results on the mean survival time $\Phi=\Theta_{2}\,\Gamma(1+\Theta_{1}^{-1})$ of the Weibull model. Panel (a) shows the distribution of $\log\Phi$ under $\mathsf{Q}_{z}^{\star}$ and Panel (b) shows the corresponding marginal IM contour for $\log\Phi$.

### v3

@@ 3.4 Back to a possibilistic IM
The previous subsections described a probabilistic approximation $\mathsf{Q}_{z}^{\star}$ of the possibilistic IM, which could be sampled with relative ease. Relevant summaries could be readily approximated using the samples from $\mathsf{Q}_{z}^{\star}$, and subsequently used to approximate the IM’s summaries. The challenge is that, for each summary, there are mathematical limits to how well a probabilistic approximation can match the target possibilistic one; for some summaries, the approximation is accurate while for others it’s not. The question is if one can convert the probabilistic approximation $\mathsf{Q}_{z}^{\star}$ to a possibilistic approximation that better represents the target IM while retaining the computational efficiency.
One simple way forward is to note that the probabilistic approximation just described can be interpreted as a possibility-to-probability transform, which has an inverse probability-to-possibility transform. Let $r_{z}:\mathbb{T}\to\mathbb{R}$ be a (possibly data-dependent) ranking function on the parameter space, where larger values of $r_{z}(\theta)$ indicate that $\theta$ is “higher ranked” in some meaningful sense. Good examples of ranking functions are, first, the density function $q_{z}^{\star}$ corresponding to the distribution $\mathsf{Q}_{z}^{\star}$ and, second, the likelihood function $L_{z}$ of the underlying model. With both of these choices, the meaningfulness of the ranking is clear. Taking the ranking function to be the density $q_{z}^{\star}$ of $\mathsf{Q}_{z}^{\star}$ gives the tightest contours, similar to how highest density sets have smallest volume among those with a fixed probability content, but “tightest contours” isn’t the objective. Moreover, there may be computational challenges associated with use of the density-based ranking. After all, $q_{z}^{\star}$ would need to be estimated using, say, kernel methods applied to the samples from $\mathsf{Q}_{z}^{\star}$ as described above, which is non-trivial when $d>1$.
@@
$\hat{\pi}_{z}(\theta)=\mathsf{Q}_{z}^{\star}\bigl{\{}r_{z}(\Theta)\leq r_{z}(\theta)\bigr{\}},\quad\theta\in\mathbb{T}.$ (13)
@@
To see that $\sup_{\theta}\hat{\pi}_{z}(\theta)=1$ and, hence, that this transformation defines a genuine possibility contour, just take $\theta$ equal (or converging) to a “highest ranked” value according to $r_{z}$. In the examples that follow, the approximation in (13) is called a stitched IM contour.
@@
Computationally, the contour $\hat{\pi}_{z}$ in (13) is far more efficient than that in (6). The reason being that the Monte Carlo samples in the former are fixed whereas, in the latter, the samples generally depend on the $\theta$ at which the contour is being evaluated. But is $\hat{\pi}_{z}$ is a good approximation of $\pi_{z}$? Towards this, consider the genuine inner probabilistic approximation $\mathsf{Q}_{z}^{\text{inn}}$ of $\Pi_{z}$ as described in Section 3.1. Then it’s easy to see that
$\mathsf{Q}_{z}^{\text{inn}}\{\pi_{z}(\Theta)\leq\pi_{z}(\theta)\}=\pi_{z}(\theta),\quad\theta\in\mathbb{T},$
where the left-hand side above is the probability-to-possibility transform of $\mathsf{Q}_{z}^{\text{inn}}$ with ranking function $r_{z}=\pi_{z}$. So, with a proper choice of ranking function, it’s theoretically possible to recover the original IM from this. The downside is that, of course, if $\pi_{z}$ was available to serve as a ranking function, then there’d be no need for any of these approximations. The same argument can be made with $\mathsf{Q}_{z}^{\text{inn}}$ and $\pi_{z}$ in the above display replaced by $\mathsf{Q}_{z}^{\star}$ and $\pi_{z}^{\text{out}}$, respectively, but the again the issue is that $\pi_{z}^{\text{out}}$ isn’t readily available. So, it’s unrealistic to expect that a stitched IM can exactly recover the original IM, but large-sample theory and my own experience suggest that using the likelihood-based ranking $r_{z}=L_{z}$ in (13) is a good standard choice.
3.5 Handling nuisance parameters
It’s often the case in applications that the full model parameter $\Theta$ is not the quantity of interest. That is, one is primarily interested in some feature $\Phi=f(\Theta)$ of the full model parameter; for example, in analysis of variance, we often care mainly about the treatment effect parameters and not so much about the error variance. Martin, 2023b presents two different strategies for marginalization in the possibilistic IM framework: the extension-based strategy is computationally and conceptually simple but generally inefficient whereas the profile-based strategy involves some problem-specific tailoring but is generally more efficient. Appendix F provides some details about these two marginalization strategies. Importantly, both strategies guarantee validity, so the more efficient profiling strategy is recommended. But use of the proposed inner probabilistic approximations adds a new twist to the marginalization story which I’ll discuss below.
The probabilistic approximation described above suggests an obvious marginalization strategy, namely, mapping the sample $\{\Theta_{m}:m=1,\ldots,M\}$ from $\mathsf{Q}_{z}^{\star}$ to a corresponding sample $\{\Phi_{m}:m=1,\ldots,M\}$ from the marginal distribution $\mathsf{Q}_{z}^{\star f}$, where $\Phi_{m}=f(\Theta_{m})$. From there, one could apply a probability-to-possibility transformation like in Section 3.4. This will be referred to as an indirect approach, in contrast to the direct approach described below. While this indirect approximation is natural and simple to carry out, a relevant question is if it’s approximating a valid marginal IM. The answer to this question depends on the mapping $f$. Martin, (2025, Theorem 6) shows that if $f$ is boundary-preserving—i.e., the boundary of $f\{C_{\alpha}(z)\}$ is the image under $f$ of the boundary of $C_{\alpha}(z)$ for each $z$ and $\alpha$, like for linear functions $f$—then $\mathsf{Q}_{z}^{\star f}$ is an inner probabilistic approximation of the extension-based marginal possibilistic IM for $\Phi$. In such cases, the indirect approach described at the start of this paragraph provides an accurate approximation of the valid extension-based marginal IM for $\Phi$.
The indirect approach need not be a bad solution when the mapping isn’t boundary-preserving since, e.g., a non-boundary-preserving map $f$ might be approximately linear in the region where $\mathsf{Q}_{z}^{\star}$ is concentrated. In general, however, one can’t guarantee that this leads to an approximation of a valid marginal IM. The point is that probabilistic marginalization through a non-linear function tends to be too concentrated and the resulting inferences are over-confident, i.e., confidence regions are too narrow, and no probability-to-possibility transform can correct this.
A safer and generally more efficient marginalization strategy is available, but requires special tailoring to the specific feature $\Phi=f(\Theta)$ of interest. First, construct the marginal IM possibilistic IM using either the extension- or profiling-based strategies described in Appendix F; of course, since profiling is more efficient, that would be my recommended strategy. Second, construct the corresponding inner probabilistic approximation to the marginal possibilistic IM. This is what I refer to as the direct approach—“direct” in the sense that the approximation is applied to the marginal IM for the specific quantity of interest. The point is that the indirect approach, which relies on probabilistic marginalization, is easy but not guaranteed to be reliable; to ensure reliability, the direct approach instead carries out possibilistic marginalization, which is provably reliable, and then applies the probabilistic approximation to the marginal IM.
4 Low-dimensional illustrations
I’ll start with some relatively simple low-dimensional examples where the proposed approximation can be visualized and, at least in some cases, the naive computation in (6) can be carried out and the proposed approximation’s accuracy can be visually assessed. A higher-dimensional example is presented in Section 6 below.
@@ 4 Low-dimensional illustrations
Let $Z=(Z_{1},\ldots,Z_{n})$ denote an iid sample from a gamma distribution with unknown parameter $\Theta=(\Theta_{1},\Theta_{2})$, where $\Theta_{1}>0$ and $\Theta_{2}>0$ are the unknown shape and scale parameters, respectively. It is straightforward to get the naive approximation (6) at any particular value of the parameter, but prohibitively expensive to carry this out over a sufficiently fine grid that spans the plausible pairs $(\theta_{1},\theta_{2})$. So, there’s a need for more computationally efficient approximation methods, and the Monte Carlo strategy of Section 3 above fits the bill. For this illustration, I’ll work with the data presented in Example 3 of Fraser et al., (1997), which consists of the survival time (in weeks) for $n=20$ rats exposed to a certain amount of radiation, modeled by a gamma distribution with unknown shape and scale parameters. Figure 2 summarizes the 5000 samples of $(\log\Theta_{1},\log\Theta_{2})$ from the proposed probabilistic approximation $\mathsf{Q}_{z}^{\star}$ developed in Section 3. This reveals that the distribution is at least approximately Gaussian. Figure 3(a) shows the exact joint contour function for $(\Theta_{1},\Theta_{2})$ along with two stitched IM approximations: one takes the ranking function to be a bivariate Gaussian density function and the other takes it to be the gamma likelihood function. Notice that the likelihood-ranking-based approximation almost perfectly matches the exact contour. For comparison, the exact contour requires more than 10 minutes of computation time, while the new Monte Carlo approximation is completed in just a few seconds.
A practically relevant and surprisingly challenging follow-up question concerns inference on the mean $\Phi=\Theta_{1}\Theta_{2}$ of the gamma distribution. IM solutions that offer exactly valid inference on $\Phi$ are presented in Martin and Liu, 2015c and Martin, 2023b , but the relevant computations can be a burden. Section 3.5 describes two approaches—direct and indirect—for carrying out this marginalization, and Figure 3(b) shows the two corresponding contours for $\Phi$, along with the exact profile-based marginal IM contour. Clearly, both the direct and indirect approaches are quite accurate in this case and, indeed, the direct approach is more accurate overall, with the indirect approach’s contour being a bit too narrow compared to the exact contour. This narrowness is inevitable because the marginal distribution for the mean $\Phi$ derived from $\mathsf{Q}_{z}^{\star}$ is not an inner probabilistic approximation of a valid marginal IM. But what the indirect approach might lack in accuracy, it makes up for in simplicity.
@@
Figure 2: Summary of the samples $(\log\Theta_{1},\log\Theta_{2})$ from $\mathsf{Q}_{z}^{\star}$ in Example 1.
@@
where $F(u)=(1+e^{-u})^{-1}$ is the logistic distribution function. The corresponding likelihood cannot be maximized in closed-form, but this is easy to do numerically, and the maximum likelihood estimator and the corresponding observed information matrix lead to the asymptotically valid inference reported by standard statistical software. For exact inference, however, the computational burden is heavier: evaluating the exact IM contour over a sufficiently fine grid of $\theta$ values is prohibitively expensive. As an alternative, the Monte Carlo sampling method presented in Section 3 is easy to implement and runs in a matter of seconds. Figure 4(a) shows the 5000 Monte Carlo samples of $(\Theta_{1},\Theta_{2})$ from $\mathsf{Q}_{z}^{\star}$ along with the stitched IM contour based on a Gaussian density ranking, closely agreeing with the approximations in Cella and Martin, (2025) and Martin and Williams, (2025).
@@
Figure 4: Summary of the results in Example 2. Panel (a) shows the samples of $(\Theta_{1},\Theta_{2})$ from $\mathsf{Q}_{z}^{\star}$ and the corresponding approximate possibility contour. Panel (b) shows the approximate possibility contour for the median lethal dose $\Lambda$.
Example 3.
@@
Following the review in Section 2 and details in Appendix F, given that $\mathsf{G}$ is a nuisance parameter, the natural strategy is to work with a relative profile likelihood,
@@
$\check{\pi}_{z}(\theta)=\mathsf{P}_{\theta,\widehat{G}}\bigl{\{}R^{\text{\sc pr}}(Z,\theta)\leq R^{\text{\sc pr}}(z,\theta)\bigr{\}},\quad\theta\in\mathbb{T}.$ (15)
Evaluation of the right-hand side via Monte Carlo boils down to sampling censoring levels from $\widehat{G}$, sampling concentration levels from $\mathsf{P}_{\theta}$, and then constructing new data sets according to (14). While this procedure is conceptually simple, naive implementation over a sufficiently fine grid of $\theta$ values is very expensive. Fortunately, the proposed Monte Carlo strategy can be readily applied to sample from a probabilistic approximation of $\check{\pi}_{z}$, from which approximately valid inference on $\Theta$ can be obtained. While only first-order asymptotic validity of this IM could be established in Cahoon and Martin, (2021), the empirical results presented there are strikingly accurate.
@@
Figure 5: Plot of samples $(\Theta_{1},\Theta_{2})$ of the Weibull shape and scale parameters from the inner probabilistic approximation $\mathsf{Q}_{z}^{\star}$ along with two corresponding joint contours: Gaussian density ranking (solid) and the asymptotic Gaussian possibility contour (dashed).
While the Weibull model is common in applications, the Weibull model parameters themselves are difficult to interpret. A relevant and interpretable feature is the mean, which is given by $\Phi=\Theta_{2}\,\Gamma(1+\Theta_{1}^{-1})$, where $\Gamma$ is the usual gamma function. Like in the gamma and logistic regression examples above, it is straightforward to obtain an approximate marginal IM for the relevant feature $\Phi$ or, equivalently $\log\Phi$, based on the samples of $(\Theta_{1},\Theta_{2})$ from $\mathsf{Q}_{z}^{\star}$. Figure 6(a) shows a histogram of the samples of $\log\Phi$ and, since there’s a slight sign of asymmetry, I use the kernel density estimate-based ranking function to construct the marginal contour for $\log\Phi$ in Figure 6(b). From here, a nominal 90% confidence interval for $\log\Phi$ can be immediately read off, i.e., $(6.41,7.74)$. Exponentiating the endpoints gives a 90% confidence interval for $\Phi$.
@@
Figure 6: Results on the mean survival time $\Phi=\Theta_{2}\,\Gamma(1+\Theta_{1}^{-1})$ of the censored Weibull model. Panel (a) shows the distribution of $\log\Phi$ under $\mathsf{Q}_{z}^{\star}$ and Panel (b) shows the corresponding marginal IM contour for $\log\Phi$.
5 Simulated-data comparisons
Numerous examples in which the IM solution compares favorably to existing methods can be found in the literature; see, e.g., Cheng et al., (2014), Martin and Lin, (2016), Cella and Martin, 2022b , and the references cited in Sections 1, 2, and 4. But the proposed approximation is at least nominally different than the IM solutions that have been previously investigated, so it’s of interest to assess how it compares to Bayesian and likelihood-based solutions. Since what’s being proposed here is a general approximation, I want to focus my comparisons on “standard” models for which the general proposal is expected to work well off-the-shelf. For the sake of brevity, I’ll focus on those “standard” models already introduced in Section 4, namely, the gamma and censored Weibull models; I’ll revisit the logistic regression example in Section 6. Since these are “standard” models, the corresponding standard Bayesian and likelihood-based solutions will perform reasonably well, so the (exact or approximate) IM solutions can’t perform significantly better. The take-away message here is that, for these two “standard” examples, the proposed approximate IM solutions perform as well—if not slightly better—than the existing Bayesian and likelihood-based solutions, which themselves are very good. Follow-up papers will consider more complex settings with model-specific improvements to the general approximation proposed here and will conduct more thorough comparisons.
Example 1 (cont).
For the gamma model, I compare the approximate IM solution to the Bayesian solution with Jeffreys prior and the likelihood-based solution (i.e., first-order asymptotic normality of the maximum likelihood estimator) in terms of coverage probability of the joint credible/confidence sets for $\Theta=(\Theta_{1},\Theta_{2})$, the gamma shape and scale. In this simulation study, I take 1000 independent samples of size $n=20$ with true $\Theta_{1}=7$ and $\Theta_{2}=3$. Figure 7(a) shows the coverage probability of the three solutions as a function of the nominal/target confidence level. All three solutions have actual coverage that nearly matches the target coverage, with both the IM and likelihood-based solutions being a bit closer to the target than the Bayes solution over a range of practical levels.
Example 3 (cont).
For the censored Weibull model, again I compare the approximate IM solution to the Bayesian solution with Jeffreys prior and the first-order likelihood-based solution in terms of coverage probability of credible/confidence sets for $\Theta=(\Theta_{1},\Theta_{2})$, the Weibull shape and scale. I take 1000 samples of size $n=20$ from the Weibull distribution with $\Theta_{1}=2$ and $\Theta_{2}=3$, but subject to Type I right censoring with respect to the ${\sf Unif}(0,4)$ distribution. Naturally, the censoring creates and additional challenge in that the IM solution described in Example 3—which relies on a plug-in estimate of the censoring distribution—is only asymptotically valid. So, it’s not clear what to expect of the proposed approximation of an only approximately valid IM. Figure 7(b) shows the coverage probability of the three solutions as a function of the nominal/target confidence level. Again, all three solutions are high-quality, but my proposed IM approximation has coverage probability at or slightly above the target level across virtually the entire range.
(a) Gamma model
(b) Censored Weibull model
Figure 7: Actual coverage probability versus the nominal/target level for three different solutions: Jeffreys prior Bayes (gray), first-order likelihood (dashed), and the proposed approximation to the possibilistic IM (solid). Diagonal dotted line is for reference.
6 A higher-dimensional illustration
Finally, I present a higher-dimensional version of the logistic regression problem in Example 2. Specifically, I consider the Pima Indians diabetes data set, widely used as a benchmark for binary classification, where the goal is to relate incidence of diabetes in Pima Indian women to eight other concomitant variables, namely, number of pregnancies, glucose concentration, blood pressure, skin thickness, insulin, BMI, diabetes pedigree function, and age. The original data set contains 768 cases. There are, however, a number of missing values, as indicated in the PimaIndiansDiabetes2 version (see the R package mlbench Leisch and Dimitriadou, 2024), and, after removing those, I’m left with $n=392$ complete cases. Compared to the previous examples, this one is relatively high-dimensional: parameter space $\mathbb{T}$ is of dimension $d=9$.
The proposed possibilistic IM approximation can easily handle problems of this magnitude, whereas such an example would be completely out of practical reach using the naive computational strategy advocated for in previous IM papers. A plot summarizing the nine-dimensional inner probabilistic approximation $\mathsf{Q}_{z}^{\star}$ is shown in Appendix G but here, for ease of visualization, Figure 8 shows two marginal IM contours: one for the linear predictor $\Phi=\Theta^{\top}x_{\text{new}}$ and one for the corresponding diabetes probability $\Psi=\text{logit}^{-1}(\Phi)$, where $x_{\text{new}}$ is the vector corresponding to the componentwise average of the eight features; note that this summary is based on the indirect marginalization strategy, utilizing the appropriate transformations of the samples of $\Theta$ from the original inner probabilistic approximation $\mathsf{Q}_{z}^{\star}$. The triangles in the two plots denote the likelihood-based asymptotic confidence limits, which are both valid and efficient in relatively large samples like available here. So, the fact that the IM’s confidence limits exactly agree with these is a sign that (a) the proposed approximation is accurate and (b) that there’s no loss of reliability here despite my use of the simple indirect marginalization strategy.
(a) Marginal contour for $\Phi=\Theta^{\top}x_{\text{new}}$
(b) Marginal contour for $\Psi=\text{logit}^{-1}(\Phi)$
Figure 8: Marginal contour for the linear predictor $\Phi=\Theta^{\top}x_{\text{new}}$ and for the corresponding probability $\Psi=\text{logit}^{-1}(\Phi)$ of a patient with feature vector $x_{\text{new}}$ having diabetes. Triangles denote the 95% likelihood-based asymptotic confidence limits.